An arrow is fired by an archer at a speed of 5m/s. Its path has the equation of 220st t=−meters after t seconds.
How high does the arrow go and find the speed covered by the rocket when its 3m above the ground?
1
Expert's answer
2020-03-02T10:21:01-0500

From the conservation of energy:


"mgH=0.5mv^2"

Height of an arrow:

"H=\frac{v^2}{2g}=\frac{5^2}{2(9.8)}=1.3\ m"



The speed covered by the rocket when its 3m above the ground:

"V(3)=\sqrt{V^2-2gh}"

"V(3)=\sqrt{220^2-2(9.8)(3)}=219.87\frac{m}{s}"